OpenStudy (anonymous):

solve by using elimination 3r-5s=-34 5r+3s=34

OpenStudy (anonymous):

\[9r-15s=-102\]\[-25r-15s=-170\] If you subtract the bottom from the top, you are left with: \[34r=68\] so r=2. By plugging 2 in for r in the first equation: \[6-5s=-34\]\[-5s=-40\]\[s=8\]

OpenStudy (anonymous):

ok where i get confused is how do I know which number to mult. each equation by

OpenStudy (anonymous):

I got r=2 but I got stuck on s

OpenStudy (anonymous):

so it would be (2,8)

OpenStudy (anonymous):

Yes, all you have to do after getting one of the variables is substitute it into one of the given equations.

OpenStudy (anonymous):

thanks for the help